### v1

# MARBLE-Net: Learning to Localize in Multipath Environment with Adaptive Rainbow Beams

Abstract
Integrated sensing and communication (ISAC) systems demand precise and efficient target localization, a task challenged by rich multipath propagation in complex wireless environments. This paper introduces MARBLE-Net (Multipath-Aware Rainbow Beam Learning Network), a deep learning framework that jointly optimizes the analog beamforming parameters of a frequency-dependent rainbow beam and a neural localization network for high-accuracy position estimation. By treating the phase-shifter (PS) and true-time-delay (TTD) parameters as learnable weights, the system adaptively refines its sensing beam to exploit environment-specific multipath characteristics. A structured multi-stage training strategy is proposed to ensure stable convergence and effective end-to-end optimization. Simulation results show that MARBLE-Net outperforms both a fixed-beam deep learning baseline (RaiNet) and a traditional k-nearest neighbors (k-NN) method, reducing localization error by more than 50% in a multipath-rich scene. Moreover, the results reveal a nuanced interaction with multipath propagation: while confined uni-directional multipath degrades accuracy, structured and directional multipath can be effectively exploited to achieve performance surpassing even line-of-sight (LoS) conditions.
I Introduction
The evolution toward 6G networks features the deep integration of sensing and communication, giving rise to the paradigm of Integrated Sensing and Communication (ISAC) [Gonzalez-Prelcic2025Six]. Leveraging large bandwidths and massive antenna arrays, ISAC enables high-precision localization as a native network service [Jabeen2025Localization]. However, this integration exposes a dichotomy in how wireless systems interact with multipath propagation. For communications, multipath is a valuable resource that enables diversity and multiplexing gains [Alkhateeb2014MIMO]; for monostatic sensing, it is often regarded as interference that causes ambiguities and degrades performance [jia2024multipath, lu2022degrees].
This challenge has spurred a paradigm shift from multipath mitigation to multipath exploitation, where multipath components are reinterpreted as rich sources of information. In multipath-rich environments, echoes via non-line-of-sight (NLoS) paths provide additional observation perspectives and spatial diversity, improving sensing accuracy, particularly for targets with fluctuating radar cross-sections (RCS) [zhang2024multipath, lv2025target]. For extended targets such as vehicles, multipath echoes can even exceed direct reflections and carry more precise motion information [zhao2025sensing]. The goal is thus to transform multipath “from enemy into friend” by exploiting its embedded geometric cues for localization tasks [liu2025multipath].
A highly efficient physical-layer mechanism for spatial exploration in wideband ISAC systems is frequency-dependent beamforming, which forms “rainbow beams” through controlled beam squint[Yan2019Wideband, luo2023beam, Cui2023NearField, gao2023integrated, Zhou_Gui_JSAC25, Zheng2025Near]. By employing true-time-delay (TTD) devices, often as part of a phase-time array (PTA) [ratnam2022joint, Mo_Jianhua_Asilomar24, cai2025hybrid, Liang2025CFARNet, nam2025joint], this technique uniquely maps different frequency subcarriers to distinct spatial locations, enabling a wide angular sector to be scanned within a single OFDM symbol. Such “single-scan” operation greatly reduces sensing overhead and latency, making it ideal for real-time applications. The beam squint effect, usually viewed as a communication impairment, is thus harnessed to encode angular information directly in the frequency domain.
Despite this efficiency, interpreting the complex multipath-rich echoes from rainbow beams remains challenging [mokri2024custom]. The superposition of numerous propagation paths forms a unique frequency-domain “fingerprint” for each target, too intricate for conventional estimators that rely on dominant LoS paths. This complexity, however, creates an opportunity for data-driven methods. Deep learning (DL) excels at modeling non-linear mappings from high-dimensional signals to geometric outputs [papageorgiou2021deep, chen2024sdoa]. While prior works apply DL for DoA estimation [chen2024sdoa] or near-field localization with fixed rainbow beams [klus2025deep], jointly leveraging multipath propagation and adaptive beam squint within a unified neural framework remains largely unexplored.
Most existing methods, including DL-based ones, rely on fixed, pre-designed rainbow beams that are non-adaptive and often suboptimal for specific environments [klus2025deep]. To overcome this limitation, we adopt a co-design paradigm that jointly optimizes the sensing beam and estimation network. By treating the beamforming parameters—phase-shifter (PS) and TTD values—as learnable weights in an end-to-end framework, the system autonomously shapes its rainbow beams and exploits environmental multipath information. Building on this idea, we propose a novel deep learning framework for single-target localization that jointly optimizes the parameters of a TTD-based rainbow beamformer and a subsequent convolutional neural network (CNN) for position estimation. Our key contributions are summarized as follows:
  • Synergistic Co-design for Adaptive Sensing: We propose a fully differentiable, end-to-end model, named MARBLE-Net (Multipath-Aware Rainbow Beam LEarning Network), that jointly optimizes a learnable rainbow beamformer and a localization CNN. This co-design enables the system to adapt its physical-layer sensing beam to the specific propagation environment via gradient-based optimization.
  • Structured Training Methodology: We propose and validate a multi-stage training strategy, which first pre-trains the prediction network with a fixed beamformer, then adapts the beamformer’s physical parameters, and concludes with an end-to-end fine-tuning. This structured approach is critical for stable convergence and superior performance. Extensive simulations show that MARBLE-Net achieves lower localization error compared with fixed-beam deep learning and traditional k-NN methods.
II System and Channel Model
We consider an uplink localization scenario where a base station (BS) aims to determine the 2D position of a single target, modeled as an unmanned aerial vehicle (UAV). The BS is equipped with a uniform linear array (ULA) of $N_{r}$ antenna elements. Both the BS and the UAV are assumed to be operating at a height of $25\text{\,}\mathrm{m}$ above the ground plane.
Figure 1: System architecture for uplink UAV localization with a phase-time array at the BS.
II-A Rainbow Beamforming
In wideband mmWave systems, the beam squint effect arises when a conventional, frequency-independent phase-shifter array is used. The physical propagation delay across the array aperture causes the beam direction to naturally shift with frequency. This is often seen as a performance-degrading impairment for communication, but can be deliberately controlled and exploited for sensing.
By augmenting the phase-shifter array with TTD elements, it becomes possible to synthesize a desired frequency-to-space mapping. The beamforming weight vector $\mathbf{w}_{m}\in\mathbb{C}^{N_{r}\times 1}$ for the $m$-th subcarrier is:
$\mathbf{w}_{m}]_{n}=\frac{1}{\sqrt{N_{t}}}\exp\left(-\mathrm{j}\left(\phi_{n}+2\pi(f_{m}-f_{0})\tau_{n}\right)\right),$ (1)
where $f_{m}$ is the frequency of the $m$-th subcarrier and $f_{0}$ is the carrier frequency, $\phi_{n}$ and $\tau_{n}$ are the phase shift and time delay applied at the $n$-th antenna. In our framework, the phase shift and time delay values are learnable parameters. They are initialized based on the strategy in [luo2023beam] to span a predefined angular range but are subsequently optimized during training.
The received uplink signal $y_{m}$ at the BS on subcarrier $m$ after receive beamforming is expressed as:
$y_{m}=\sqrt{\frac{P_{t}}{M}}\mathbf{w}_{m}^{H}\mathbf{h}_{m}x_{m}+\eta_{m},$ (2)
where $\mathbf{h}_{m}\in\mathbb{C}^{N_{r}\times 1}$ is the channel vector, $P_{t}$ is the transmit power of UE, $x_{m}$ is the transmitted symbol, and $\eta_{m}$ is i.i.d. complex additive white Gaussian noise (AWGN). The collection of received samples, $\mathbf{y}=[y_{1},\dots,y_{M}]^{T}\in\mathbb{C}^{M\times 1}$, forms the observation vector.
(a) ‘circle’ scene
(b) ‘rounded-L’ scene
(c) ‘L’ scene
Figure 2: Three multipath scenes modeled by Blender.
II-B High-Fidelity Ray-Tracing Channel Generation
To ensure our model is trained on physically realistic channel data, we employ a meticulous simulation pipeline based on the Sionna ray-tracing library. We define four distinct 3D environments to evaluate performance under varying degrees of multipath propagation:
  1. 1.
    LoS-only Scene: An unobstructed environment with no obstacles, where propagation is dominated by the direct LoS path. This serves as a baseline reference.
  2. 2.
    ‘circle’ Scene: An environment designed to generate uni-directional multipath with low angular spread. As shown in Fig. 2(a), it consists of a semicircular wall (radius $300\text{\,}\mathrm{m}$, height $50\text{\,}\mathrm{m}$), which confines most reflected paths to a narrow azimuth range. Although these paths may differ in elevation due to wall height and target position, they predominantly arrive from the same horizontal direction, forming a characteristic uni-directional multipath pattern.
  3. 3.
    ‘rounded-L’ Scene: An environment designed to generate multipath with high angular spread. As shown in Fig. 2(b), it consists of a $300\text{\,}\mathrm{m}$ radius quarter-circle arc contiguous with a $400\text{\,}\mathrm{m}$ straight wall (both $50\text{\,}\mathrm{m}$ high). This structure introduces reflective paths from more diverse directions compared to the ‘circle’ scene.
  4. 4.
    ‘L’ scene: An environment creating a rich multipath profile including higher-order reflections. It is composed of two perpendicular walls (North-South wall at $205\text{\,}\mathrm{m}$ from origin, East-West wall at $180\text{\,}\mathrm{m}$ from origin, both $50\text{\,}\mathrm{m}$ high). This geometry is representative of urban street-corner scenarios, generating both single- and double-bounce reflection paths.
For all scenes, the radio material properties are assigned to be concrete for the ground and brick for the walls. The Blender scene is imported into Sionna, and for each of the 100 000100\,000 random UAV positions, Sionna’s path solver traces interactions (reflections, refractions) to compute all valid propagation paths and synthesize the final frequency-domain channel matrix $\mathbf{H}$.
III Proposed Localization Framework
We introduce MARBLE-Net, a fully differentiable, end-to-end architecture that maps the raw channel matrix to a 2D coordinate estimate. It comprises two cascaded modules: a learnable rainbow beamformer ($\mathcal{M}_{1}$) and a prediction network ($\mathcal{M}_{2}$). As illustrated in Fig. 3, the channel data are first processed by $\mathcal{M}_{1}$ to generate a power spectrum, which is subsequently fed into $\mathcal{M}_{2}$ for localization. The prediction error is then backpropagated through both modules, enabling joint optimization of the physical-layer beamforming and the prediction network parameters.
Figure 3: Data flow of the MARBLE-Net architecture.
III-A Learnable Rainbow Beamformer ($\mathcal{M}_{1}$)
This module acts as the physical-layer front end. Its trainable parameters consist solely of the PS and TTD vectors, $\bm{\phi},\bm{\tau}\in\mathbb{R}^{N_{r}}$. Given the channel matrix, it computes the received spectrum $\mathbf{y}$ using (1) and (2), and produces the log-magnitude power spectrum $\mathbf{P}_{\text{dB}}\in\mathbb{R}^{M}$ with each element defined as $\mathbf{P}_{\text{dB},m}=10\log_{10}(|y_{m}|^{2}+\epsilon)$. This logarithmic transformation effectively compresses the wide dynamic range of wireless signals, facilitating stable learning.
III-B Prediction Network ($\mathcal{M}_{2}$)
The prediction network is a 1D CNN designed to regress the target’s 2D Cartesian coordinates from the 1D power spectrum.
  • Feature Extractor: A stack of five Conv1d layers with progressive downsampling extracts hierarchical spectral features, effectively capturing multi-scale patterns induced by multipath interference.
  • Regression Head: An MLP with three hidden layers processes the flattened feature vector. The final output layer has 2 neurons for the $(x,y)$ coordinates.
The output is rescaled by the maximum possible distance to yield the final estimate. The entire MARBLE-Net model is trained to minimize the MSE loss.
III-C Multi-Stage Training Strategy
Optimizing a coupled system of a physical-layer model and a neural network presents unique challenges due to the complex, non-convex loss landscape. To ensure stable convergence and achieve superior performance, we propose a structured multi-stage training strategy, akin to curriculum learning. This approach methodically optimizes different parts of the MARBLE-Net.
  • Stage 1: Prediction Network Pre-training. Train only the prediction network ($\mathcal{M}_{2}$) while keeping the beamformer ($\mathcal{M}_{1}$) fixed with its initial parameters. This phase allows the network to learn a mapping from a non-adaptive, but structured, input spectrum to the target’s location, establishing a strong feature extraction baseline.
  • Stage 2: Beamformer Adaptation. Freeze the weights of the prediction network ($\mathcal{M}_{2}$) and train only the learnable PS and TTD parameters of the beamformer ($\mathcal{M}_{1}$). In this stage, the beamformer is forced to adapt its physical sensing pattern to generate a power spectrum that is more “interpretable” for the pre-trained network, thereby minimizing the downstream localization error.
  • Stage 3: End-to-End Fine-tuning. Unfreeze all parameters and jointly train the entire MARBLE-Net model. This allows for fine-grained adjustments between the beamformer and the prediction network, leading to a synergistic optimization of the end-to-end system.
IV Simulation Results
IV-A Simulation Setup
We generate a dataset of 100 000100\,000 samples for each of the four scenes using the Sionna pipeline, with key parameters listed in Table I. Each dataset is split into training (80%), validation (10%), and testing (10%) sets. All models are implemented in PyTorch and trained using the Adam optimizer. MARBLE-Net is trained using the multi-stage strategy described in Section III-C.
TABLE I: Key Simulation Parameters
Parameter Symbol Value
Start Frequency $f_{0}$ $28\text{\,}\mathrm{GHz}$
Bandwidth $W$ $\approx$380\text{\,}\mathrm{MHz}$\$
Subcarrier Spacing $\Delta f$ $240\text{\,}\mathrm{kHz}$
Number of Subcarriers $M$ 1584
BS Antennas (ULA) $N_{r}$ 128
UAV Antenna $N_{t}$ 1
UAV Transmit Power $P_{t}$ 23, 18, 13 dBm
BS/UAV Height - $25\text{\,}\mathrm{m}$
UAV-BS Distance Range - $5\text{\,}\mathrm{m}$, $200\text{\,}\mathrm{m}$]
UAV Azimuth Range - $^{\circ}, 6060^{\circ}$]
Ray-Tracing Max Depth - 2 (LoS, Reflection)
Dataset Size - 100 000100\,000 samples
TABLE II: Performance comparison and ablation study of different localization methods across the four scenes
Scene Metric k-NN RaiNet [klus2025deep] RaiNet MARBLE-Net MARBLE-Net
($K=5$) (fixed) (adaptive) (fixed) (adaptive)
LoS-only Localization RMSE (m) 15.73 4.02 3.62 3.94 3.46
Angle RMSE () 0.13 1.21 1.11 1.55 1.17
Range RMSE (m) 15.73 3.82 3.44 3.84 3.16
‘circle’ Localization RMSE (m) 16.63 5.34 4.94 4.77 4.28
Angle RMSE () 0.18 1.61 1.47 1.36 5.49
Range RMSE (m) 16.63 5.15 4.78 4.65 4.06
‘rounded-L’ Localization RMSE (m) 13.32 4.18 3.75 3.93 2.96
Angle RMSE () 0.17 1.25 1.20 1.51 1.41
Range RMSE (m) 13.32 3.99 3.51 3.71 2.78
‘L’ Localization RMSE (m) 5.85 1.88 1.55 1.06 0.88
Angle RMSE () 0.14 0.71 0.54 0.43 0.35
Range RMSE (m) 5.84 1.63 1.38 0.94 0.81
(a) LoS-only Scene
(b) ‘circle’ Scene
(c) ‘rounded-L’ Scene
(d) ‘L’ Scene
Figure 4: Spatial distribution of the localization error of the MARBLE-Net (adaptive) at $P_{t}=$23\text{\,}\mathrm{d}\mathrm{B}\mathrm{m}$\$ across the four scenes.
IV-B Baseline
IV-B1 k-Nearest Neighbors (k-NN)
To benchmark against a traditional, non-learning-based method, we implement a k-NN algorithm with $K=5$. A codebook is constructed by storing the received power spectra (from a fixed initial beam) for 80,000 training samples (codewords) and their corresponding locations. At inference time, the spectrum from a test target is compared against all codebook entries using Euclidean distance. The coordinates of the $K$ nearest neighbors are averaged to produce the location estimate.
IV-B2 RaiNet
As a point of comparison, the baseline RaiNet given in [klus2025deep] adopted a fixed rainbow beam and only trained the prediction network. This represents a conventional DL-based approach with a non-adaptive sensing beam. The RaiNet architecture consists of a feature extractor and a classifier. The feature extractor contains three 1D convolutional layers, followed by a classifier with three fully connected layers. The Tanh function is used as the primary activation function throughout the network.
IV-C Performance Comparison and Ablation Study
The localization performance of the proposed MARBLE-Net, alongside several baselines, is comprehensively evaluated across the four distinct environments. The results, summarized in Table II, provide a detailed ablation study on the impact of multipath complexity, joint beamformer-network training, and network architecture.
First, it is evident that all deep learning approaches dramatically outperform the traditional k-NN baseline. In the ‘circle’ scene, for instance, k-NN yields a large localization RMSE of $16.63\text{\,}\mathrm{m}$, underscoring its inability to handle the complex signal fingerprints created by multipath propagation. Template-matching methods like k-NN fail when the direct relationship between the signal spectrum and geometric location is severely distorted, rendering the pre-compiled codebook ineffective.
To deconstruct the sources of performance gain, we compare four DL-based configurations: RaiNet with a fixed beamformer (‘RaiNet (fixed)’), RaiNet with multi-stage training (‘RaiNet (adaptive)’), MARBLE-Net’s architecture with a fixed beamformer (‘MARBLE-Net (fixed)’), and the fully optimized MARBLE-Net. This allows us to isolate the effects of network architecture and joint training.
The Benefit of Joint Training: A key finding is the critical importance of joint end-to-end training. Comparing the ‘fixed’ and ‘multi-stage’ trained results for both architectures reveals a consistent performance improvement. For RaiNet in the ‘circle’ scene, multi-stage training reduces the RMSE from $5.34\text{\,}\mathrm{m}$ to $4.94\text{\,}\mathrm{m}$, a relative improvement of 7.5%. However, the benefit is far more pronounced for MARBLE-Net, where the RMSE drops from $4.77\text{\,}\mathrm{m}$ to $4.28\text{\,}\mathrm{m}$, a substantial 10.3% improvement. This demonstrates that while a simple network like RaiNet gains moderately from adapting the beamformer, the more complex architecture of MARBLE-Net synergizes much more effectively with the adaptive physical layer, unlocking noticeable performance gains.
The Nuanced Impact of Multipath: The results in Table II reveal a complex, non-monotonic relationship between multipath complexity and localization accuracy. When moving from the ‘LoS-only’ scene (MARBLE-Net RMSE of $3.46\text{\,}\mathrm{m}$) to the highly complex ‘circle’ scene, performance for all models degrades. For MARBLE-Net, the RMSE increases to $4.28\text{\,}\mathrm{m}$. This suggests that the dense, uni-directional multipath in the ‘circle’ scene introduces ambiguity that is difficult to resolve, even for an adaptive system.
However, the ‘rounded-L’ scene tells a different story. Counter-intuitively, MARBLE-Net’s performance in this moderately complex scene (RMSE of $2.96\text{\,}\mathrm{m}$) is superior to both the ‘LoS-only’ and ‘circle’ scenes. This strongly indicates that MARBLE-Net is not merely mitigating multipath but is actively exploiting the structured, directional NLoS paths provided by the ‘rounded-L’ scene’s geometry. These paths act as a source of valuable spatial diversity, allowing the model to refine its position estimate. In contrast, the fixed-beam models fail to leverage this, showing worse performance in the ‘rounded-L’ scene than in the LoS scene, confirming that adaptation is the key to turning specific multipath from an impairment into a resource.
This trend of multipath exploitation is further exemplified by the ‘L’ scene. In this environment with two perpendicular walls, the localization error for MARBLE-Net drops to $0.88\text{\,}\mathrm{m}$. This is a significant improvement over all other scenarios, including LoS-only ($3.46\text{\,}\mathrm{m}$). This result suggests that the simple, planar reflections from the ‘L’ scene provide highly stable and geometrically informative NLoS paths, which the adaptive beamformer and network learn to use as ‘virtual anchors’, dramatically refining the position estimate.
The Benefit of Network Complexity: Finally, the role of network architecture is highlighted. With its fixed beamformer, MARBLE-Net’s more complex architecture (‘MARBLE-Net (fixed)’) offers a consistent but marginal advantage over the simpler ‘RaiNet (fixed)’ across all scenes. This suggests that simply increasing the number of network parameters is insufficient. It is the combination of a high-capacity prediction network and a co-adapted, learnable physical-layer beamformer that enables the system to learn the intricate mapping from multipath-rich channel observations to precise location estimates, unlocking the large performance leap seen in the ‘MARBLE-Net (adaptive)’ results.
Figure 5: CDF of localization error for MARBLE-Net (adaptive) at $P_{t}=$23\text{\,}\mathrm{d}\mathrm{B}\mathrm{m}$\$ across all scenes.
IV-D Qualitative and Error Distribution Analysis
To gain deeper insight into the system’s behavior, we visualize the spatial error distribution of the MARBLE-Net (adaptive) model in Fig. 4. The visualization aligns perfectly with the quantitative results in Table II. The ‘L’ scene (Fig. 4d) shows the lowest error, with the vast majority of the area in dark blue (low error). The ‘rounded-L’ scene (Fig. 4c) also shows excellent performance, appearing visually superior to the ‘LoS-only’ scene (Fig. 4a), which has larger green and yellow high-error regions. In stark contrast, the ‘circle’ scene (Fig. 4b) exhibits large swathes of high-error (yellow, orange, and red) zones. These high-error zones correspond to areas where the uni-directional multipath effects are most severe, creating signal ambiguities that even the adaptive MARBLE-Net struggles to resolve completely. This visualization clearly illustrates that structured multipath (‘L’, ‘rounded-L’) is exploited, while uni-directional multipath (‘circle’) remains a significant challenge.
The cumulative distribution function (CDF) of localization error, shown in Fig. 5, offers additional insight into performance consistency across environments. The ‘L’ scene exhibits the most favorable distribution, with the curve rising steeply and reaching about 95% probability at a $2\text{\,}\mathrm{m}$ error threshold. The ‘rounded-L’ scene also performs well, clearly outperforming the ‘LoS-only’ case. In contrast, the ‘circle’ scene shows a long-tailed distribution due to unidirectional ambiguous multipath, with large errors not fully displayed in the figure. These observations are consistent with Table II, confirming that structured multipath is beneficial, whereas uni-directional multipath remains detrimental.
IV-E Impact of Transmit Power
Figure 6: Localization RMSE vs. UAV Transmit Power for MARBLE-Net (adaptive).
We evaluate the system’s robustness to noise by varying the UAV transmit power, as shown in Fig. 6. As expected, the localization RMSE improves for all scenes with increased transmit power due to the higher SNR. Notably, the performance hierarchy across scenes remains consistent, with the ‘L’ and ‘rounded-L’ scenes outperforming the ‘LoS-only’ case, and the ‘circle’ scene performing the worst. This confirms that the geometric impact of multipath (whether beneficial or detrimental) is consistent across different noise levels.
IV-F Complexity and Latency Analysis
TABLE III: Complexity and latency comparison
Method Parameters / Memory Avg. Latency/Sample
MARBLE-Net 13.96 M params $0.0172\text{\,}\mathrm{m}\mathrm{s}$
RaiNet [klus2025deep] 1.95 M params $0.0180\text{\,}\mathrm{m}\mathrm{s}$
k-NN (GPU) 484 MB codebook $2.2122\text{\,}\mathrm{m}\mathrm{s}$
Table III compares computational complexity and latency, measured on an NVIDIA RTX 5880 GPU with dual AMD EPYC 7Y43 CPUs. Although MARBLE-Net has far more parameters (13.96 M vs. 1.95 M for RaiNet), its average inference latency ($0.0172\text{\,}\mathrm{m}\mathrm{s}$) remains comparable to RaiNet ($0.0180\text{\,}\mathrm{m}\mathrm{s}$). Both models achieve over two orders of magnitude faster inference than the k-NN baseline ($2.21\text{\,}\mathrm{m}\mathrm{s}$), which depends on a 484 MB codebook and costly neighbor search. These results confirm the proposed model’s balance between accuracy and efficiency, enabling real-time deployment.
V Conclusion
This paper has introduced MARBLE-Net, an end-to-end deep learning framework that jointly optimizes an adaptive rainbow beamformer and a localization network for ISAC systems. By leveraging ray-tracing-generated datasets across environments of varying complexity, we have demonstrated that a data-driven approach can improve localization performance. Our results show that co-designing the sensing beam with the neural network processor, guided by a structured multi-stage training strategy, allows the system to learn to exploit multipath information, reducing localization error from $5.85\text{\,}\mathrm{m}$ and $1.88\text{\,}\mathrm{m}$ to $0.88\text{\,}\mathrm{m}$ in a multipath-rich scene. This work validates that for high-precision localization in complex environments, the joint design of the physical-layer sensing strategy and the higher-layer processing network is not merely beneficial, but essential. Our findings also highlight a nuanced interaction with the propagation environment: while confined uni-directional multipath can hinder performance, we show that structured and directional multipath can be actively exploited to achieve accuracy surpassing even LoS-only conditions.
Future work will extend this framework to 3D and multi-target localization, and explore online adaptation in dynamic environments.